# PDFMathTranslate: Scientific Document Translation Preserving Layouts

### v2

@@ 1 Introduction
The advances in science and technologies spread and evolve in different languages  Von Gizycki ; Montgomery . Yet, language barriers hinder the diffusion and development of scientific progress Ramírez-Castañeda (2020); Hwang (2005); Ammon (2012). For example, while 98% publications in science are written in English Liu ; Ammon , the language only has 7.3% native speakers and no more than 20% speakers worldwide Bahji et al. . Consequently, the majority of human beings are hindered by language barriers from the advances in science and technology. To overcome the barrier, there are attempts such as those from international organizations UNESCO and academic efforts in improving Machine Translation (MT) over text Brown et al. (a); Vaswani et al. (2017); Zhu et al. ; Johnson et al. ; Sennrich et al. .
However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012). In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures has rich and important meanings. Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
@@ 2 Architecture and Design
PDFMathTranslate is designed to translate documents while preserving their original layouts. First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A). Next, it detects the layouts to extract the layouts and textual contents from the document. Third, the texts are translated using the selected translation service, such as GPT-4, DeepL, Google, or Ollama (see Appendix A). Finally, the translated text and previously detected layouts are re-rendered as a translated document with preserved layouts.
Technically, we design the architecture for precision, flexibility, and efficiency. We introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, which we will explain in the following three subsections.
Precise layout parser
Preserving layouts requires precise layout extraction. To precisely parse the position of document elements, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering. In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx. The model is another version of a SOTA solution in object detection, Yolov10 Wang et al. , and is fast and accurate in the specific task Zhao et al. .
To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version. The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardware.
Flexible translation middleware
Translation middleware offers flexibility in terms of language diversity, supported services, and customization capabilities. One, PDFMathTranslate supports at least 56 languages (see Appendix A), including popular ones such as English, and some languages shared by relatively smaller communities, such as Cherokee. Users can freely select any of these languages as either the input or output language when translating documents.
@@
Three, PDFMathTranslate provides additional customizability in translation. We designed a customization feature prompting strategies that potentially increase the quality of translation or its adaptability in domain-specific tasks. This feature allows users to exploit advanced prompting strategies such as few-shots Brown et al. (b), Chain-of-Thought Wei et al. , Role-playing Shanahan et al. , etc.
Efficient streaming flow
@@ 6 Sustainability
Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar . To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
First, we increase developer exposure by prominently featuring contributors on the project’s homepage via third-party services and regularly highlighting recent contributions in project documentation (e.g., 1970-01-01 / Supports Google translation (by @author_handle). This strategy has successfully incentivized 44 global developers who collectively contributed tens of thousands of lines of code.
@@ 8 Acknowledgments
The authors thank all 44 contributors [^5] They are 7shi, Byaidu, Copilot, Cycloctane, Hanaasagi, IuvenisSapiens, JEFF-dev-ui, Tql-ws1, Wybxc, YadominJinta, Zxis233, alohays, aseaday, awwaawwa, borcation, charles7668, chiu0602, czz404, damaoooo, dependabot[bot], domonnss, eltociear, hellofinch, highkay, hotwa, imClumsyPanda, kharkover, kidach1, lintian233, mydreamworldpolly, namazuchin, qqueing, reycn, tastelikefeet, timelic, treeleaves30760, tylzh97, ws051682, wx-11, xxnuo, xyzyx233, yidasanqian, ymattw, zqqian; ordered alphabetically, retrieved through the official GitHub API by March 2025) of the open-sourced project on GitHub; platforms allowing us to host demos (Hugging Face, ModelScope); services we used in the demo (Google Translate); and all comments from users.
@@ 9 Ethics and Broader Impacts
Our work has already made a significant positive impact within both multilingual scientific and open-source communities. However, as a flexible, widely used tool, it raises potential ethical concerns regarding the copyright of documents. Translating documents without proper permissions could challenge the intellectual property rights of scientific works and innovations. We welcome suggestions from experts in intellectual property to mitigate the concern.
@@ A.1 Officially supported language and services
Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 1
models Ollama Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
AzureOpenAI, Zhipu, ModelScope, Silicon, Gemini, Azure,
@@
Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 2
output Portuguese (Brazil), Bulgarian, Catalan, Cherokee, Croatian,
languages Czech, Danish, Dutch, English (US), Estonian, Filipino,
Finnish, French, German, Greek, Gujarati, Hebrew, Hindi,
Hungarian, Icelandic, Indonesian, Italian, Japanese,
Kannada, Korean, Latvian, Lithuanian, Malay, Malayalam, Marathi,
Norwegian, Polish, Portuguese (Portugal), Romanian, Russian,
Serbian, Chinese (Simplified), Slovak, Slovenian, Spanish, Swahili,
Swedish, Tamil, Telugu, Thai, Chinese (Traditional), Turkish,
Urdu, Ukrainian, Vietnamese, Welsh
@@
    2 Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.

### v4

@@ 1 Introduction
Science and technologies diffuse and develop in different languages  Von Gizycki (1973); Montgomery (2013). Yet, language barriers hinder the diffusion and development of scientific progress Ramírez-Castañeda (2020); Hwang (2005); Ammon (2012). For example, while 98% publications in science are written in English Liu ; Ammon (2012), the language only has 7.3% native speakers and no more than 20% speakers worldwide Bahji et al. (2023). Consequently, the majority of human beings are hindered by language barriers from the advances in science and technologies. To overcome the barrier, there are the attempts from international organizations for inclusive science  [^1] See https://www.unesco.org/en/open-science and the attempts from academia in improving Machine Translation (MT) over text Brown et al. (1990); Vaswani et al. (2017); Zhu et al. (2020); Johnson et al. (2017); Sennrich et al. (2016).
However, text-based machine translation fails to address the unique challenges posed by the layouts in technical translation Schubert (2012) where the elaborate access structure matters (pp. 351-352). Non-textual elements in scientific and technical documents are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures have rich and important meanings. Thus, text-based machine translation ignoring such important information appear insufficient to address the barrier in scientific and technical document translation, hindering the progress of science, technologies, and society.
@@ 2 Architecture and Design
PDFMathTranslate is designed to translate documents while preserving their original layouts. First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service. Next, it detects the layouts to extract the layouts and textual contents from the document. Third, the texts are translated using the selected translation service, such as GPT-4, DeepL, Google, or Ollama (see Appendix A). Finally, the translated text and previously detected layouts are re-rendered as a translated document with preserved layouts.
Technically, we design the architecture for precision, flexibility, and efficiency. These three principles are realized by a precise parser, a flexible translation middle-ware, and an efficient workflow, detailed in following subsections.
2.1 Precise layout parser
Preserving layouts requires precise layout extraction. To precisely parse the position of document elements, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering. In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx. The model is another version of a SOTA solution in object detection, Yolov10 Wang et al. (2024), and is fast and accurate in the specific task Zhao et al. (2024).
To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX)  [^2] See https://onnx.ai/ standard and a pytorch version. The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardware.
2.2 Flexible translation middleware
Translation middleware offers flexibility in terms of language diversity, supported services, and customization capabilities. One, PDFMathTranslate supports at least 194 languages (see Appendix A), including popular ones such as English, and some languages shared by relatively smaller communities, such as Cherokee. Users can freely select any of these languages as either the input or output language when translating documents.
@@ 2.2 Flexible translation middleware
Three, PDFMathTranslate provides additional customizability in translation. We designed a customization feature prompting strategies that potentially increase the quality of translation or its adaptability in domain-specific tasks. This feature allows users to exploit advanced prompting strategies such as few-shots Brown et al. (2020), Chain-of-Thought Wei et al. (2022), Role-playing Kong et al. (2024); Shanahan et al. (2023), etc.
2.3 Efficient streaming flow
@@ 6 Sustainability
Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar (2010). To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
First, we increase developer exposure by prominently featuring contributors on the project’s homepage and regularly highlighting recent contributions (e.g., 1970-01-01 / Supports Google translation (by @author_handle). This strategy has successfully incentivized 44 global developers who collectively contributed tens of thousands of lines of code.
@@ 8 Acknowledgments
The authors thank all 44 contributors [^7] They are 7shi, Byaidu, Copilot, Cycloctane, Hanaasagi, IuvenisSapiens, JEFF-dev-ui, Tql-ws1, Wybxc, YadominJinta, Zxis233, alohays, aseaday, awwaawwa, borcation, charles7668, chiu0602, czz404, damaoooo, dependabot[bot], domonnss, eltociear, hellofinch, highkay, hotwa, imClumsyPanda, kharkover, kidach1, lintian233, mydreamworldpolly, namazuchin, qqueing, reycn, tastelikefeet, timelic, treeleaves30760, tylzh97, ws051682, wx-11, xxnuo, xyzyx233, yidasanqian, ymattw, zqqian; ordered alphabetically, retrieved through the official GitHub API by March 2025) of the open-sourced project on GitHub; platforms allowing us to host demos (Hugging Face, ModelScope); services we used in the demo (e.g., Google Translate); redeem codes from Immersive Translate; public API access from SiliconFlow; and all comments from users.
@@ 9 Ethics and Broader Impacts
Our work has made a significant positive impact within both multilingual scientific and open-source communities. However, as a flexible, widely used tool, it raises potential ethical concerns regarding the copyright of documents. Translating documents without proper permissions could challenge the intellectual property rights of scientific works and innovations. We welcome suggestions from experts in intellectual property to mitigate the concern.
@@ A.1 Officially supported language and services
In Table 2, we enumerate the officially supported services, models, and languages. However, the actual number may be significantly higher due to the extensibility of services and models.
@@
Service or Google (default), 302.AI, Bing, DeepL, DeepLX, Ollama, Ali Qwen, 23 1
models Ollama, Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
AzureOpenAI, Zhipu, ModelScope, Silicon Cloud, Gemini, Azure,
@@
Input and Abkhaz, Acehnese, Acholi, Afrikaans, Albanian, Alur, Amharic, 194 2
output Arabic, Armenian, Assamese, Awadhi, Aymara, Azerbaijani, Balinese,
languages Bambara, Bashkir, Basque, Batak Karo, Batak Simalungun,
Batak Toba, Belarusian, Bemba, Bengali, Betawi, Bhojpuri,
Bikol, Bosnian, Breton, Bulgarian, Buryat, Cantonese,
Catalan, Cebuano, Chichewa (Nyanja), Chinese (Simplified),
Chinese (Traditional), Chuvash, Corsican, Crimean Tatar,
Croatian, Czech, Danish, Dinka, Divehi, Dogri, Dombe,
Dutch, Dzongkha, English, Esperanto, Estonian, Ewe,
Fijian, Filipino (Tagalog), Finnish, French,
French (French), French (Canadian), Frisian,
Fulfulde, Ga, Galician, Ganda (Luganda), Georgian,
German, Greek, Guarani, Gujarati, Haitian Creole,
Hakha Chin, Hausa, Hawaiian, Hebrew, Hiligaynon,
Hindi, Hmong, Hungarian, Hunsrik, Icelandic, Igbo,
Iloko, Indonesian, Irish, Italian, Japanese, Javanese,
Kannada, Kapampangan, Kazakh, Khmer, Kiga, Kinyarwanda,
Kituba, Konkani, Korean, Krio, Kurdish (Kurmanji),
Kurdish (Sorani), Kyrgyz, Lao, Latgalian, Latin, etc.
@@
    2 Based on the default service (Google Translate) and interface (CLI).